OpenStudy (anonymous):

To find the height of a pole, a surveyor moves 160 feet away from the base of the pole and then, with a transit 6 feet tall, measures the angle of elevation to the top of the pole to be 62 degrees. What is the height of the pole?

OpenStudy (anonymous):

Can you try to draw it?

OpenStudy (anonymous):

@zepdrix @dan815 @TheSmartOne

OpenStudy (anonymous):



OpenStudy (anonymous):

Cool, so do you know the trig functions?

OpenStudy (anonymous):

tangent

OpenStudy (anonymous):

I just don't really know how to enter it on a Calculator

OpenStudy (anonymous):

Okay, what is the equation for tangent? Do you know it?

OpenStudy (anonymous):

no, i need help

OpenStudy (anonymous):

wait I think, but im not sure

OpenStudy (anonymous):

Okay, well the equation is: \[ \tan(\text{angle}) = \frac{\text{opposite}}{\text{adjacent}} \]So we have: \[ \tan(62^\circ) = \frac{x}{160\ \text{ft}} \]So now, we must solve for \(x\).

OpenStudy (anonymous):

haha I know this i put that in the drawing but i ment the steps after this

OpenStudy (anonymous):

How do you isolate \(x\)?

OpenStudy (anonymous):

Where are you stuck? Is your calculator in degree mode?

OpenStudy (anonymous):

I can get down all the way to these steps. I don't know how I would enter it on the calculator though

OpenStudy (anonymous):

First of all\[ x = \tan(62^\circ) \times 160\ \text{ft} \]On you calculator, is should have a \(\tan\) button. You should see what happens when you press it.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

what would the final answer be after i enter it @wio

OpenStudy (anonymous):

oh ok i see now, thanks you so much for your help <3